Wednesday, February 7, 2018

PFS Upgrade Series, Day 14: Getting Into the Groove

This is part of a series of posts about upgrading an instrument at Las Campanas Observatory. If you want to start at the beginning, it's here.

Last night was our third night on sky, and it really started to feel like observing with the "old" PFS. The last two weeks, I/we have been super focused installing new components and getting them to work, and then getting the instrument as a whole working the way we want and expect it to. That has been tiring but also extremely satisfying in its own right -- for instance, when we finally figured out how to do the optical alignment with the HASO, or when we (Jeff really) got the slit mechanism working. But it wasn't until last night that it seemed to all come together, and come back to the scientific goal of the whole project, finding and characterizing planets orbiting other stars. The data we are taking on this run may or may not actually be used to do this planet-finding work directly, but it most certainly will help us figure out how to take and reduce data in the near future to do just that. I think it's rare for astronomers to be so fully connected to a project as I feel I am with this one, and I'm not even dealing closely with the software side of things, which is a *huge* part! I feel very, very lucky to have this opportunity to work with this team of folks on this instrument. 

Before diving into 1x1 binning observations of some stable stars, we did another quick focus test of the instrument, as well as played around with the voltages on the CCD to mitigate the weird bias pattern I mentioned yesterday. In the photo below, Steve is keeping track of what voltages Christoph inputs into the CCD controller and what the images look like; we wanted to see if saturating the detector had different effects with the different voltages. Turns out it didn't, so we settled on a setting of A_HIGH=7 and A_LOW=-6, although we (or at least Christoph and me) don't know exactly what those values mean. But it evens out the bias structure the best! We still have those residual bright columns on the right side, but we hope those can be masked out in the reduction. Our contact at STA,  Kasey, said, "Based on the shape I think we are seeing residual P+ from the thinning process. I don't think we will be able to reduce this more without jeopardizing other aspects of the device.  The column defect is now generating more charge so it is near the surface." 

We still like to take paper logs. Call us old-fashioned. 

The weather last night was pretty good, but the seeing deteriorated during the middle of the night to >1''! Yick. It settled down towards the end of the night, but we were almost through by then. 

Clay seeing is in red, Baade (the other Magellan telescope) in blue. See that big spike around 5:00 UT? 

We also got a few errors from the GUI last night, after trying to do loops of exposures, so I'll have to ask Christoph about that. He added a nice feature in the GUI where we can reconnect to the CCD controller, rather than having to go manually reboot it, but it still takes a minute or two, which we'd like to avoid wasting on non-engineering nights. Tonight the plan is to observe the same stars we've been observing every night (ha, does anyone remember Pinky and the Brain?), and transfer the data to Paul Butler at Carnegie DTM in Washington, DC for more in-depth analysis. We sent him the first two nights of data already, although I don't know that he knows it...I shall email him now! 

No comments:

Post a Comment